OpenStudy (anonymous):

solve the equation for the indicated variable \[A=1/2h(B+b) for \] h

OpenStudy (diyadiya):

\[A=\frac{1}{2}h(B+b) \] is this the equation ?

OpenStudy (anonymous):

yes, i couldnt figure out how to make he h in the middle

OpenStudy (diyadiya):

Ok so first multiply both sides by 2

OpenStudy (diyadiya):

\[2 \times A = 2 \times \frac{1}{2}h(B+b)\]

OpenStudy (diyadiya):

Now what happens to the 2's in the right side?

OpenStudy (anonymous):

do they cancel out?

OpenStudy (diyadiya):

Yes!!

OpenStudy (diyadiya):

Now what do you have ?

OpenStudy (anonymous):

2xA=h(B+b) ?

OpenStudy (diyadiya):

Right!! :)

OpenStudy (diyadiya):

Now you have to isolate h what do you think you can do ?

OpenStudy (anonymous):

do you subtract the h from both sides?

OpenStudy (diyadiya):

Nope its multipplication so you cant subtract it we divide both sides by (B+b)

OpenStudy (diyadiya):

2A = h x (B+b) so divide both sides by (B+b) so that we can isolate h (since (B+b) will cancell out )

OpenStudy (anonymous):

2xA/B+b the other cancells out correct?

OpenStudy (diyadiya):

yes!!! :)\[\Huge h= \frac{2A}{(B+b)}\]

OpenStudy (anonymous):

just 2A i didnt need the multiply sign anymore got it

OpenStudy (diyadiya):

:)

OpenStudy (anonymous):

ty, that made it easy your the best!

OpenStudy (diyadiya):

You're Welcome :)) Glad to help!!!